OpenStudy (anonymous):

If a rock falls from a height of 38 meters on Earth, the height H (in meters) after x seconds is approximately H(x)=38-4.9x^2. When is the height of the rock 11 meters? The answer is 2.35 seconds, but I don't see how they got that answer. Can someone please explain this to me?

OpenStudy (stamp):

\[H(x)=38-4.9x^2\]Find x for H(x)=11 m.\[11=38-4.9x^2,\ 27=4.9x^2\]\[x=\sqrt{\frac{27}{4.9}}=2.35s\]

OpenStudy (anonymous):

Thanks! :)